# Reinforcement Learning for LLM Reasoning Under Memory Constraints

### v1

@@ IV-D Discussion
V Limitations
Our study has several limitations. First, we do not compare against full model fine-tuning baselines, as this was infeasible given our single 40GB GPU constraint. It remains an open question whether full fine-tuning would mitigate some of the challenges we observed with group-level credit assignment.
Second, while we benchmarked a GRPO baseline adapted for LoRA, we did not perform extensive hyperparameter sweeps. GRPO’s reliance on gradient accumulation through more forward and backward passes increases wall-clock time, making robust tuning impractical in our setup. As a result, we cannot conclusively claim that our methods will outperform GRPO in more optimized regimes.
Finally, our experiments focus on a single base model (Qwen2-1.5B). Although this choice enables controlled comparisons, future work should validate the generality of our findings across a broader range of model architectures and scales.
VI Future Work
A promising direction for future research is to evaluate T-SPMO on tasks involving longer and more semantically diverse chains of thought, such as those found in the benchmark MATH. In these settings, surface-level token matching may be insufficient for effective credit assignment. We hypothesize that augmenting the prefix-matching mechanism in T-SPMO with a semantic-aware structure—such as a fuzzy-matching prefix tree based on latent representations or paraphrase clustering—could improve the generalizibility of token-level RL and reduce sensitivity to lexical variation.
As for S-GRPO, next steps could include variations on the sampling algorithm for tokens that contribute to loss. While we found that sampling with a strong bias towards the beginning of a completion is a simple and effective method, exploring more sophisticated algorithms would likely be a promising research direction.
VII Conclusions
We introduced two reinforcement learning methods—Stochastic Group Relative Policy Optimization (S-GRPO) and Token-Specific Prefix Matching Optimization (T-SPMO)—designed to improve reasoning performance in large language models under constrained computational budgets. Both methods are critic-free and operate within a LoRA fine-tuning framework, making them practical for training on a single 40GB GPU. Across two benchmarks, SVAMP and multi-digit multiplication, our approaches significantly outperform zero-shot baselines, with T-SPMO achieving particularly strong results on arithmetic tasks.

### v2

@@ 4.4 Discussion
5 Conclusions
We presented two reinforcement learning methods—Stochastic Group Relative Policy Optimization (S-GRPO) and Token-Specific Prefix Matching Optimization (T-SPMO)—designed specifically for LoRA-based fine-tuning of large language models under tight memory and compute constraints. Both methods are critic-free and intentionally reduce the number of output tokens contributing to the gradient, enabling efficient and stable training. By aligning the algorithmic structure with the limitations of low-rank adaptation, our approaches demonstrate that substantial reasoning gains are possible without full-sequence loss computation or heavyweight model updates. Across two benchmarks, SVAMP and multi-digit multiplication, our approaches significantly outperform zero-shot baselines, with T-SPMO achieving particularly strong results on arithmetic tasks.